OpenStudy (anonymous):

at the school carnival annika's class earned 120 dollars in 1 hour and 30 minutes.at thiss rate how much will they earn during the 6 hours the carnival is opened

OpenStudy (anonymous):

anyone i need help

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@harpreetsk

OpenStudy (anonymous):

ie in one min they earn 120/90 in 6hours they will earn 120X60X6/90=43200/90=480

OpenStudy (texaschic101):

120/1.5 = x/6 -- 120 in 1.5 hrs = x in 6 hrs cross multiply (1.5)(x) = (120)(6) 1.5x = 720 x = 720/1.5 x = 480

OpenStudy (anonymous):

lol @shelbygt520

OpenStudy (anonymous):

mhmmm most annoying person I've ever dealt with... hell he might be worse than my little brother

OpenStudy (anonymous):

little brother haaa lol